1 The Irish language

The file irish.dtx^{1} defines all the language definition macros for the Irish language.

For this language currently no special definitions are needed or available.

The macro \LdfInit takes care of preventing that this file is loaded more than once, checking the category code of the @ sign, etc.

1 ⟨∗code⟩

 2 \LdfInit{irish}\captionsirish

When this file is read as an option, i.e. by the \usepackage command, irish could be an ‘unknown’ language in which case we have to make it known. So we check for the existence of \l@irish to see whether we have to do something here.

 3 \ifx\l@irish\@undefined
 4 \@nopatterns{irish}
 5 \adddialect\l@irish0\fi

The next step consists of defining commands to switch to (and from) the Irish language.

\captionsirish The macro \captionsirish defines all strings used in the four standard documentclasses provided with L^{A}TEX.

 6 \addto\captionsirish{%
 7 \def\prefacename{Preface}%    <-- needs translation
 8 \def\refname{Tagairt\'{\i}}%
 9 \def\abstractname{Achoimre}%
10 \def\bibname{Leabharliosta}%
11 \def\chaptername{Caibidil}%
12 \def\appendixname{Aguis\'{\i}n}%
13 \def\contentsname{Cl\'ar \'Abhair}%
14 \def\listfigurename{L\'ear\'aid\'{\i}}%
15 \def\listtablename{T\'abla\'{\i}}%
16 \def\indexname{Inn\'eacs}%
17 \def\figurename{L\'ear\'aid}%
18 \def\tablename{T\'abla}%
19 \def\partname{Cuid}%
20 \def\enclname{faoi iamh}%
21 \def\ccname{cc}%                 abrv. `c\'oip chuig'
22 \def\headtoname{Go}%
23 \def\pagename{Leathanach}%
24 \def\seename{see}%    <-- needs translation
25 \def\alsoname{see also}%    <-- needs translation
26 \def\proofname{Proof}%    <-- needs translation

27 }

\dateirish The macro \dateirish redefines the command \today to produce Irish dates.

28 \def\dateirish{%
29 \number\day\space \ifcase\month\or
30 Ean\'air\or Feabhra\or M\'arta\or Aibre\'an\or
31 Bealtaine\or Meitheamh\or I\'uil\or L\'unasa\or
32 Me\'an F\'omhair\or Deireadh F\'omhair\or

1The file described in this section has version number v1.0e and was last revised on 1996/12/23. A contribution was made by Marion Gunn.

33 M\'{\i} na Samhna\or M\'{\i} na Nollag\fi
34 \space \number\year}

\extrasirish \noextrasirish The macro \extrasirish will perform all the extra definitions needed for the Irish language. The macro \noextrasirish is used to cancel the actions of \extrasirish. For the moment these macros are empty but they are defined for compatibility with the other language definition files.

35 \addto\extrasirish{}
36 \addto\noextrasirish{}

The macro \ldf@finish takes care of looking for a configuration file, setting the main language to be switched on at \begin{document} and resetting the category code of @ to its original value.

37 \ldf@finish{irish}

38 ⟨/code⟩